# Generalized Hypercube Queuing Models with Overlapping Service Regions

### v2

@@ 2 Proposed framework
This section develops a framework for assessing districting plans with overlapping patrol units for service systems. To gain insight into the strategy of overlapping patrol, we consider a moving-server queuing system with multiple servers. Typically, each server (e.g., a patrol service unit) is assigned to one particular service region, where the server has primary responsibility. When a server is not busy serving any active calls, it traverses its home region to perform preventive patrol. A key feature of our model is that multiple servers are allowed to be jointly responsible for a third overlapping service region besides their respective home regions, as illustrated in Figure 5. For the reader’s convenience, Table 1 contains summary definitions of notations frequently used in the paper.
Table 1: Summary of key notations
Section Notation Description
2.1 $I$ Total number of servers.
=$\mathcal{I}=\{i=1,\dots,I\}$ Set of indices of servers as well as their primary service regions.
=$\mathcal{E}=\{e\in\mathcal{I}\}$ Set of overlapping service regions.
$\lambda_{i}$ Arrival rate of primary service region $i$.
$\lambda_{e}$ Arrival rate of overlapping service region $e$.
$\lambda$ Arrival rate within the entire region of interest.
$\mu_{i}$ Service rate of server $i$.
$\mu$ Total service rate of all servers.
2.2 $n_{i}\in\mathbb{Z}_{*}$ Status of server $i$. Server $i$ is idle if $n_{i}=0$ and busy if $n_{i}>0$.
$\eta_{e,u}(i)\in[0,1]$ Probability of choosing server $i$ in overlapping region $e$.
$B=(n_{i})_{i\in\mathcal{I}}$ A state (node) in the hyperlattice.
$\mathcal{U}=\{u=1,2,\dots\}$ Set of state indices ordered by the tour algorithm.
$Q=(q_{uv})_{u,v\in\mathcal{U}}$ Transition rate matrix for the hyperlattice queue.
2.3 $N_{k}$ A state in the aggregate (birth-death) model
$\bar{\lambda}_{k}$ Birth rate of aggregated state $N_{k}$.
$\bar{\mu}_{k}$ Death rate of aggregated state $N_{k}$.
$U_{K}$ Number of states where the total number of calls in the system is less than $K$.
$\mathcal{U}_{K}=\{u=1,\dots,U_{K}\}$ Set of states for the truncated hyperlattice queue.
$Q_{K}=(q_{uv})_{u,v\in\mathcal{U}_{K}}$ Transition rate matrix for the truncated hyperlattice queue.
2.4 $\rho_{i}$ Fraction of time that server $i$ is busy serving calls.
$\rho_{i,e}/\rho_{i,i}$ Fraction of dispatches that send server $i$ to each sub-region $e$ (or $i$).
@@ 2.1 Problem setup and key assumptions
In this section, we describe our problem setup and several key assumptions. Consider a service system patrolled by multiple servers, where each server is allocated to a specific area. Unlike conventional service systems with separate service territories, our approach permits servers to have overlapping service areas. The entire service system can be represented by an undirected hypergraph [6] denoted by $(\mathcal{I},\mathcal{E})$. In this hypergraph, =$\mathcal{I}=\{i=1,\dots,I\}$ denotes the set of vertices, each representing an individual server, with $I$ being the total number of servers in the system. The set $\mathcal{E}$ represents hyperedges, each of which corresponds to a group of servers denoted as $e\in\mathcal{E}\subseteq\mathcal{I}$, jointly overseeing an overlapping area. The cardinality of $\mathcal{E}$ is denoted by $E$. Figure 6 presents the hypergraph of a three-server system as an example.
@@
(a) Two-server system
(b) Three-server system
Figure 5: Examples of three service systems with overlapping patrol. The white area $\mathcal{P}_{i}$ indicates primary service region $i$ that can only be served by server $i$ and the shaded area $\mathcal{O}_{e}$ indicates overlapping service region $e=\{i,j\}$ that can be jointly served by servers $i$ and $j$.
Figure 6: The hypergraph of a three-server system, where the black dots represent vertices and the shaded areas represent hyperedges.
Assume service calls arrive in the system according to a general stationary arrival process. The arrivals of calls in primary service region $i$ and overlapping service region $e$ follow time-homogeneous Poisson processes with rate $\lambda_{i}$ and $\lambda_{e}$, respectively. Let $\lambda=\sum_{i\in\mathcal{I}}\lambda_{i}+\sum_{e\in\mathcal{E}}\lambda_{e}$ be the aggregate call arrival rate in the entire region. In addition, we assume that the service region size is small enough and the arrival rate is high enough to ensure that the servers are continuously busy, causing the queuing delay to dominate the travel time. This assumption will justify the birth-death process we will be using later.
The service process is described as follows. The server starts to process the call as soon as it arrives at the location where the call is reported. The service time of each server is independent of the location of the customer and the history of the system; The service time of server $i$ for any call for service has a negative exponential distribution with mean $1/\mu_{i}$. Let $\mu=\sum_{i\in\mathcal{I}}\mu_{i}$ denote the aggregate service rate in the entire region. After finishing the service, the server stays at the same location until receiving a new call from the dispatch center. For ergodicity, it is necessary that $\lambda<\mu$ should hold.
We assume each server has its own queue, with the dispatch policy being state-dependent. For the sake of simplicity in presenting our findings, we start our discussion with a traditional setup that many service systems typically use. Then we will show that our assumptions can be applied more broadly across various service systems. In the traditional context, we adopt the following dispatch policy: If a call arrives in the service region of a server and finds the server available (randomly choose a server if there is more than one), then it goes into service upon the server’s arrival. When there is no server available, the dispatch policy can be discussed in the following three situations:
  1. 1.
    If the call arrives in the primary service region of a server, it joins the end of the queue of the server.
  2. 2.
    If a call is received in the overlapping service region, the call will be randomly assigned to one of the available servers that are in charge of this region.
  3. 3.
    If a call is received in the overlapping service region and there is no available server, the call will be randomly assigned to one of the responsible servers, and wait in its corresponding queue until the server becomes available.
An example of the dispatch policy for a two-server system with one overlapping service region is presented in Figure 7.
(a) Both idle
(b) One busy
(c) Both busy
Figure 7: Queue networks with two servers and two queues when (a) both servers are idle, (b) one of the servers is busy, and (c) both servers are busy, where $\eta$ is the probability of assigning the call to server $i$.
2.2 Hyperlattice queuing model for overlapping patrol
We propose a generalized hypercube queuing model, called the hyperlattice model, to capture the operational dynamics of systems with overlapping service regions. The system state depends on the status of all the servers in this system, and the number of calls in each queue to be processed.
@@ 2.2.1 State representation
These states can be represented by a hyperlattice in dimension $I$. Each node of the hyperlattice corresponds to a state $B=(n_{i})_{i\in\mathcal{I}}$ represented by a tuple of numbers, where non-negative integer $n_{i}\in\mathbb{Z}_{*}$ indicates the status of server $i$. Server $i$ is idle if $n_{i}=0$ and busy if $n_{i}>0$. The value of $n_{i}-1$ represents the number of calls waiting in the queue of server $i$ when the server is busy. Figure 8 (a) gives an example of the state space of a hyperlattice queuing model for a system with two servers. It is worth noting that the state space in a hyperlattice queuing model can be divided into two parts, with one part consisting of states that possess at least one available server (referred to as non-saturated states) and the other part comprising states in which all servers are busy (referred to as saturated states), as shown in Figure 9.
(a) State space
(b) Aggregated state space
(c) Case 1: for $n_{1}\geq 1$ and $n_{2}<1$
(d) Case 2: for $n_{1},n_{2}\geq 1$
(e) Case 3: for $n_{1}<1$ and $n_{2}\geq 1$
Figure 8: The state space in a hyperlattice queuing model for a system with two servers of equal service rate $\mu/2$. The state space is shown in (a) as an overview, with each state denoted by the pair $(n_{1},n_{2})$. Green arrows indicate downward transitions and red arrows indicate upward transitions. (b) presents the birth-death model by aggregating states that share the same value of $n_{1}+n_{2}$. In (c), (d), and (e), the details of the state space are depicted in three distinct scenarios.
(a) Two servers
(b) $I$ servers
Figure 9: An illustration of two types of states in hyperlattice queuing models with different numbers of servers. The area colored red signifies the saturated states that have no idle servers, whereas the empty area represents the non-saturated states that have at least one server that is unoccupied.
@@ 2.2.2 Touring algorithm
(a) State space
(b) Tour of all states with $k\leq 3$
(c) Tour within $N_{2}$
Figure 10: A hyperlattice queuing model for a system with three service regions. (a) shows part of the state space of the hyperlattice for $k=1,2,3$; (b) presents the tour of all the states with $k\leq 3$, where solid arrows represent the tour within an aggregated state and dashed arrows represent the tour across aggregated states; (c) shows the sequence of all the states in the aggregated state $N_{2}$, where the number of balls separated by dashed circles represents the status of server $i$ ($n_{i}$).
Input: Number of servers $I$; Output: The sequence of ordered states $\mathcal{B}$;
@@
       $\mathcal{B}_{k}\leftarrow\{(n_{i})\in\mathbb{Z}_{*}^{I}\mid\sum_{i\in\mathcal{I}}n_{i}=k\}$; // $|\mathcal{B}_{k}|=\binom{k+I-1}{I-1}$
@@
We simplify the tour problem by decomposing it into $K$ sub-problems. Each sub-problem $k\leq K$ first identifies the states in which exactly $k$ calls are staying in the system (either waiting in the queue or being served), and then proceeds to solve the enumerative combinatorics problem by searching for $\binom{k+I-1}{I-1}$ possible combinations in a pre-determined sequence. It is important to note that the sequence generated by each sub-problem is just one of numerous possible tours of the cutting plane in the hyperlattice. Figure 10 presents a hyperlattice with $I=3$, along with its corresponding tour of all the states. The combinatorics can be translated into the “stars and bars” [10], i.e., finding all possible positions of $I-1$ dashed circles that separate $k$ balls into $I$ groups, as illustrated in Figure 10 (c). The number of balls being separated in group $i$ can be regarded as $n_{i}$. Algorithm 1 summarizes the tour algorithm that generates the sequence $B_{0},B_{1},\dots$. For convenience, we index the states based on the order produced by the tour algorithm. The state index is represented by $u\in\mathcal{U}=\{0,1,2,\dots\}$.
@@ 2.2.3 Transition rate matrix
Proposition 1 (Upward transition rate).
@@
$\eta_{e,u}(i)$ $=0,$ $~{}\forall i\notin e,e\in\mathcal{E},$ (2)
$\sum_{i\in\mathcal{I}}\eta_{e,u}(i)$ $=1,$ $~{}\forall e\in\mathcal{E}.$ (3)
The following argument justifies the proposition. The first term of (1) suggests that, for an arbitrary server $i$, it must respond to a call received in its primary service region $i$. The second term of (1) means that the server $i$ needs to respond to a call received in the overlapping service region $e$ according to the dispatch policy $\eta_{e,u}$. It is important to emphasize that the dispatch policy $\eta_{e,u}$ is state-dependent and can be defined based on the user’s needs, offering a general framework that is apt for elucidating both random and deterministic dispatch strategies. For example, the one outlined in Section 2.1 represents a specific case of the dispatch policy, in which the assignment of server $i$ to the call received in overlapping service region $e$ can be discussed in two scenarios: (a) When all servers $i\in e$ are busy ($u:n_{i}>0,\forall~{}n_{i}\in B_{u}$), this call will join the queue of server $i$ based on the probability $\eta_{e,u}(i)$. (b) Conversely, when there is at least one server idle ($u:n_{i}=0,n_{i}\in B_{u}$), the server $i$ will be assigned to this call based on the probability $\eta_{e,u}(i)$. In Appendix A, we also demonstrate how this proposition can be distilled to represent a rudimentary system where, at most, only two servers are permitted to patrol one overlapping service region.

### v4

@@ 3 Proposed framework
This section develops a framework for assessing districting plans with overlapping patrol units for service systems. To gain insight into the strategy of overlapping patrol, we consider a moving-server queueing system with multiple servers. Typically, each server (e.g., a patrol service unit) is assigned to one particular service region, where the server has primary responsibility. When a server is not busy serving any active calls, it traverses its home region to perform preventive patrol. A key feature of our model is its consideration of a more flexible state-dependent policy, allowing servers to collaborate and jointly serve a third overlapping service region within a spatial queueing framework, as illustrated in Figure 3.1. For the reader’s convenience, Table EC1 contains summary definitions of notations frequently used in the paper.
@@ 3.1 Problem setup and key assumptions
In this section, we describe our problem setup and several key assumptions. Consider a service system patrolled by multiple servers, where each server is allocated to a specific area and has its own queue. Unlike conventional service systems with separate service territories, our approach permits servers to have overlapping service areas. The entire service system can be represented by an undirected hypergraph (Bretto 2013) denoted by $(\mathcal{I},\mathcal{E})$. In this hypergraph, $\mathcal{I}=\{i=1,\dots,I\}$ denotes the set of vertices, each representing an individual server, with $I$ being the total number of servers in the system. The set $\mathcal{E}$ represents the hyperedges, where each hyperedge $e\in\mathcal{E}$ is a subset of $\mathcal{I}$, i.e., $\mathcal{E}\subseteq 2^{\mathcal{I}}$. Each hyperedge corresponds to a group of servers jointly overseeing an overlapping area. The cardinality of $\mathcal{E}$ is denoted by $E$. Figure 3.1 presents the hypergraph of a three-server system as an example.
@@
[…]
Examples of 2- and 3-server systems with overlapping patrols. The white region $\mathcal{P}_{i}$ denotes the primary service area of server $i$, which can only be served by that server. The shaded region $\mathcal{O}_{e}$ represents the overlapping service area $e=\{i,j\}$, which can be jointly served by servers $i$ and $j$.
[…]
The hypergraph of a three-server system. The black dots represent vertices and the shaded areas represent hyperedges.
Assume service calls arrive in the system according to a Poisson process with constant arrival rate. The arrivals of calls in primary service region $i$ and overlapping service region $e$ follow time-homogeneous Poisson processes with rate $\lambda_{i}$ and $\lambda_{e}$, respectively. Let $\lambda=\sum_{i\in\mathcal{I}}\lambda_{i}+\sum_{e\in\mathcal{E}}\lambda_{e}$ be the total call arrival rate in the entire region.
The service process is described as follows. The server begins processing the call as soon as it arrives at the location where the call is reported. The service time of each server is independent of the customer’s location and the system’s history. The service time of server $i$ for any call for service has a negative exponential distribution with mean $1/\mu_{i}$. Let $\mu=\sum_{i\in\mathcal{I}}\mu_{i}$ denote the total service rate in the entire region. After completing the service, the servers return to or remain within their primary service territory, regardless of where the previous call originated. For ergodicity, it is necessary that $\lambda<\mu$ should hold (a detailed stability condition is provided in Section 3.2.5).
3.2 Hyperlattice queueing model for overlapping patrol
We propose a generalized hypercube queueing model, called the hyperlattice model, to capture the operational dynamics of systems with overlapping service regions. The system state depends on the status of all the servers in this system, and the number of calls in each queue to be processed.
@@ 3.2.1 State representation
These states can be represented by a hyperlattice in dimension $I$. Each node of the hyperlattice corresponds to a state $B=(n_{i})_{i\in\mathcal{I}}$ represented by a tuple of numbers, where a non-negative integer $n_{i}\in\mathbb{Z}_{+}$ indicates the status of server $i$. Server $i$ is idle if $n_{i}=0$ and busy if $n_{i}>0$. The value of $n_{i}-1$ represents the number of calls waiting in the queue of server $i$ when the server is busy. Figure 3.2.1 (a) gives an example of the state space of a hyperlattice queueing model for a system with two servers. It is worth noting that the state space in a hyperlattice queueing model can be divided into two parts, with one part consisting of states that possess at least one available server (referred to as non-saturated states) and the other part comprising states in which all servers are busy (referred to as saturated states), as shown in Figure 3.2.1.
[…]
The state space in a hyperlattice queueing model for a system with two servers. The state space is shown in (a) as an overview, with each state denoted by the pair $(n_{1},n_{2})$. Green arrows indicate downward transitions and red arrows indicate upward transitions. (b) presents the birth-death model by truncating states at an upper bound $K=2$. In (c), (d), and (e), the details of the state space are depicted in three distinct scenarios.
[…]
An illustration of two types of states in hyperlattice queueing models with different numbers of servers. The area colored red signifies the saturated states that have no idle servers, whereas the empty area represents the non-saturated states that have at least one server that is unoccupied.
@@ 3.2.2 Touring algorithm
[…]
A hyperlattice queueing model for a system with three service regions. (a) shows part of the state space of the hyperlattice for $k=1,2,3$; (b) presents the tour of all the states with $k\leq 3$, where solid arrows indicate transitions within states with the same number of units, while dashed arrows represent transitions between states with different numbers of units; (c) shows the sequence of all states within $k=2$, where the number of balls separated by dashed circles represents the status of server $i$ ($n_{i}$).
Input: Number of servers II; Output: The sequence of ordered states $\mathcal{B}$;
@@
 $\mathcal{B}_{k}\leftarrow\{(n_{i})\in\mathbb{Z}_{+}^{I}\mid\sum_{i\in\mathcal{I}}n_{i}=k\}$, where $|\mathcal{B}_{k}|=\binom{k+I-1}{I-1}$;
@@
We simplify the tour problem by decomposing it into $K$ sub-problems. Each sub-problem $k\leq K$ first identifies the states in which exactly $k$ calls are staying in the system (either waiting in the queue or being served), and then proceeds to solve the enumerative combinatorics problem by searching for $\binom{k+I-1}{I-1}$ possible combinations in a pre-determined sequence. It is important to note that the sequence generated by each sub-problem is just one of numerous possible tours of the cutting plane in the hyperlattice. Figure 3.2.2 presents a hyperlattice with $I=3$, along with its corresponding tour of all the states. The combinatorics can be translated into the “stars and bars” (Flajolet and Sedgewick 2009), i.e., finding all possible positions of I1I-1 dashed circles that separate $k$ balls into $I$ groups, as illustrated in Figure 3.2.2 (c). The number of balls being separated in group $i$ can be regarded as $n_{i}$. Algorithm 1 summarizes the tour algorithm that generates the sequence $B_{0},B_{1},\dots$. For convenience, we index the states based on the order produced by the tour algorithm. The state index is represented by $u\in\mathcal{U}=\{0,1,2,\dots\}$.
3.2.3 Dispatch policy
We define a class of dispatch policies within a general state-dependent framework. Let $\mathcal{H}$ denote the set of all Markovian dispatch policies. Our state-dependent dispatch policy, $\eta\in\mathcal{H}$, is defined as follows: When a call arrives at location $s\in\mathcal{S}$, we first determine whether $s$ belongs to a primary service region or an overlapping service region. If $s\in\mathcal{P}_{i}$ for some server $i\in\mathcal{I}$, the call is routed directly to server $i$. If $s$ belongs to an overlapping region, i.e., $s\in\mathcal{O}_{e}$ for some server set $e\in\mathcal{E}$, then the call can be served by any server in $e$. In this case, the dispatch policy is characterized by a discrete distribution $\eta_{e,u}=(\eta_{e,u}(i))_{i\in e}$, where $\eta_{e,u}(i)$ denotes the probability of assigning the call to server $i\in e$. The corresponding dispatch policy is denoted as $\eta=\left\{\eta_{e,u}\mid e\in\mathcal{E},u\in\mathcal{U}\right\}$. This broad class of state-dependent policies includes canonical examples such as “Join-the-Shortest-Idle-Queue” in Weng et al. (2020).
[…]
Queue networks with two servers and two queues. (a) both servers are idle, (b) one of the servers is busy, and (c) both servers are busy, where $p$ is the probability of assigning the call to server 1.
A simple example of such a dispatch policy for a two-server system with a single overlapping service region is illustrated in Figure 3.2.3, where the distribution $\eta_{e,u}$ is designed to prioritize the idle server. Specifically, in state $u$, if only one server is idle, the dispatch policy assigns the call exclusively to that server, i.e., $\eta_{e,u}(i)=1$ for the idle server $i$. Otherwise, if both servers are either available or busy, the call is assigned randomly, with $\eta_{e,u}(1)=p$ and $\eta_{e,u}(2)=1-p$ for some fixed probability $p\in[0,1]$.
@@ 3.2.4 Transition rate matrix
Proposition 3.1 (Upward transition rate)
@@
$\eta_{e,u}(i)$ $=0,$ $~\forall i\notin e,e\in\mathcal{E},u\in\mathcal{U},$ (2)
$\sum_{i\in\mathcal{I}}\eta_{e,u}(i)$ $=1,$ $~\forall e\in\mathcal{E},u\in\mathcal{U}.$ (3)
The following argument justifies the proposition. The first term of (1) suggests that, for an arbitrary server $i$, it must respond to a call received in its primary service region $i$. The second term of (1) means that the server $i$ needs to respond to a call received in the overlapping service region $e$ according to the dispatch policy $\eta_{e,u}$. It is essential to note that the dispatch policy $\eta_{e,u}$ is state-dependent and can be defined based on the user’s specific needs, providing a general framework that is suitable for elucidating both random and deterministic dispatch strategies. For example, Figure 3.2.3 represents a specific case of the dispatch policy, in which the assignment of server $i$ to the call received in overlapping service region $e$ can be discussed in two scenarios: (a) When all servers $i\in e$ are busy ($u:n_{i}>0,\forall~n_{i}\in B_{u}$), this call will join the queue of server $i$ based on the probability $\eta_{e,u}(i)$. (b) Conversely, when there is at least one server idle ($u:n_{i}=0,n_{i}\in B_{u}$), the server $i$ will be assigned to this call based on the probability $\eta_{e,u}(i)$.
3.2.5 Markov chain representation
The hyperlattice model introduced above can be formulated as a continuous-time Markov chain (CTMC) on a discrete state space. Specifically, it is characterized by the state space $B$ and the transition rate matrix $Q$. The original state space is defined as $B:=\mathbb{Z}_{+}^{|\mathcal{I}|}$. For convenience, we introduce a touring algorithm that establishes a one-to-one mapping from the high-dimensional state space $B$ to a one-dimensional index state space $\mathcal{U}:=\mathbb{Z}_{+}$. Additionally, let $e_{i}$ denote a vector with all elements zero except for the $i$-th entry, which is set to 1. The transition rate matrix $Q=(q_{uv})_{u,v\in\mathcal{U}}$ is then given by:
$q_{uv}=\begin{cases}\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e,u}(i)\lambda_{e}&\text{if }B_{v}=B_{u}+e_{i},\quad\forall i\in I,\\ \mu_{i}&\text{if }B_{v}=B_{u}-e_{i},\left(B_{u}\right)_{i}>0,\quad\forall i\in I,\\ -\lambda-\sum_{i\in I:\left(B_{u}\right)_{i}>0}\mu_{i}&\text{if }B_{v}=B_{u},\\ 0&\text{otherwise.}\end{cases}$ (4)
For the hyperlattice queueing model to be stable, i.e., to attain a steady-state distribution, the corresponding CTMC must be ergodic. To ensure ergodicity, we impose the following stability condition for general policies.
{assumption}
[…]
$\sum_{e\in\mathcal{A}}\lambda_{e}<\sum_{i\in\mathcal{I}:\,\exists\,e\in\mathcal{A},(i,e)\in\tilde{\mathcal{E}}}\mu_{i},\quad\forall\,\mathcal{A}\subseteq\mathcal{N}.$
Assumption 3.2.5 is analogous to Hall’s marriage theorem (Hall 1987). When this condition holds, by max-flow/min-cut theorem (Ford Jr and Fulkerson 1956), there exists flows $\Lambda_{ie}\in\mathbb{R}_{+}^{|\mathcal{I}|\times|\mathcal{N}|}$ with
$\sum_{i\in\mathcal{I}}\Lambda_{ie}=\lambda_{e}\quad(\forall e\in\mathcal{N}),\quad\sum_{e\in\mathcal{N}}\Lambda_{ie}\leq\mu_{i}-\delta_{i}\quad(\forall i\in\mathcal{I}),$
for some $\delta_{i}>0$ and $\Lambda_{ie}=0$ whenever $(i,e)\notin\tilde{\mathcal{E}}$. Then, we define a state-independent policy $\eta$, under which each arrival from primary or overlapping region $e\in\mathcal{N}$ is assigned to server $i\in\mathcal{I}$ according to a fixed probability distribution:
$\eta_{e}(i)=\frac{\Lambda_{ie}}{\sum_{i^{\prime}\in\mathcal{I}}\Lambda_{i^{\prime}e}},\quad\forall i\in\mathcal{I},e\in\mathcal{N}.$ (5)
Lemma 3.2
Under Assumption 3.2.5, there exists a state-independent policy $\eta$ as defined in (5) that ensures stability of the queueing system.
The proof of Lemma 3.2 follows directly from Assumption 3.2.5 and the max-flow/min-cut theorem. Note that this stability condition applies generally to state-independent policies. More sophisticated state-dependent policies (e.g., Join-the-Idle-Queue) can also be designed to ensure stability whenever Hall’s condition is satisfied.